OpenStudy (anonymous):

What are the solutions to the following system of equations? y = x2 + 3x − 7 3x − y = −2 (3, 11) and (−3, −7) (11, 3) and (−3, −7) (3, 11) and (−7, −3) No Real Solutions

OpenStudy (anonymous):

help

OpenStudy (mertsj):

Replace y in the first equation with 3x+2 from the second equation.

OpenStudy (anonymous):

you can substitute values

OpenStudy (anonymous):

is that right mertsj

OpenStudy (mertsj):

Then get it equal to 0 and try to factor it.

OpenStudy (anonymous):

What is one of the solutions to the following system of equations? y2 + x2 = 53 y − x = 5 (−10, −5) (−7, −2) (7, 2) (5, 10)

OpenStudy (anonymous):

last one

OpenStudy (mertsj):

Did you plug it into both equations to see if that choice makes both equations true?

OpenStudy (mertsj):

In the first answer choice, what is the x value?

OpenStudy (anonymous):

CAN YOU JUST give me the answer please

OpenStudy (mertsj):

How would that help you learn?

OpenStudy (anonymous):

iit homeowrk and i just want the answer so tommorow my teacher can show me how to do it again

OpenStudy (anonymous):

just substitute the values of x and y... if both sides are equal then the answer is correct.

OpenStudy (anonymous):

which is

OpenStudy (anonymous):

alright i got b

OpenStudy (anonymous):

is it y^2 + x^2 = 53 ?

OpenStudy (anonymous):

or a

OpenStudy (anonymous):

b and c is out of the picture

OpenStudy (anonymous):

i mean c and d

OpenStudy (anonymous):

(-7,-2) is correct

OpenStudy (mertsj):

What did you get?